# Not All Contexts Are Equal: Teaching LLMs Credibility-aware Generation

### v1

@@ 5 Experiments
To demonstrate the effectiveness of our framework in handling flawed information in real-world QA scenarios, we conduct comprehensive experiments under three scenarios within the CAGB. All these results verify the effectiveness of the proposed CAG framework and the corresponding training algorithm. Additionally, our models maintain robustness even with the increasing noise in the context. In the following sections, we will discuss our experiments and conclusions in detail.
@@ 5.1 Setup
We compare our method with the following three strategies incorporated with 6 LLMs across various scales:
@@
We evaluate advanced models, including ChatGPT (gpt-3.5-turbo-0613), LLaMA-2-7B, 13B, 70B, Vicuna-7B-v1.5 and Mistral-7B-Instruct (Jiang et al., 2023). Additionally, we create a dataset mirroring the model training data but without credibility annotations and with initial answers, on which we fine-tune the same base model, and named the trained model vanilla IFT.
Figure 3: The performance of LLMs under varying noise ratios, which denote the proportions of retrieved noise documents. As the noise ratio increases, the performance of previous methods significantly deteriorates; in contrast, our approach maintains stable prediction quality even in high noise ratio, attributed to its enhanced ability to identify and prioritize accurate information.
@@
We use LLaMA-2-7B, LLaMA-2-13B and Mistral-7B as our base models. To provide relevance scores, we use SPLADE (Formal et al., 2021) as our retriever. Our model training is based on the Fastchat framework and carry out on two A100-80G GPUs. For all language models, we incorporate 3-shot QA examples within the prompt. We set the temperature parameter to 0.01 during inference. We employ Exact Match (EM) (Rajpurkar et al., 2016) as the primary evaluation metric for all datasets. The prompts used for our evaluation are provided in the Appendix A.3.
@@ 5.2 Overall Results
The main results of the three scenarios are presented in the Table 2, we can clearly see that our model efficiently understands and utilizes credibility information to provide more accurate and credible responses. In the following, we analyze the experimental results in detail:
@@
In scenarios including open-domain QA, time-sensitive QA, and misinformation pollutedQA, existing LLMs, including ChatGPT and LLaMA-2-70B, face challenges due to interference from flawed information. In the retrieval-based open-domain QA scenario, the average EM score for ChatGPT is only 41.5%, while the EM score for LLaMA-2-70B is 44.1%. All models exhibit low performance on the Musique, NewsPollutedQA, which are characterized by high ratios of flawed information. The method of reranking using externally provided relevance scores can assist the model to a certain extent, as the model is sensitive to the order of documents (Xie et al., 2023; BehnamGhader et al., 2023).
@@
The models, developed through training on LLaMA 7B, 13B, and Mistral 7B with CAG, not only exhibit improved reliability in its outputs but also excel in new, challenging situations, including time-sensitive QA and misinformation polluted QA. This performance, achieved within an open-domain QA framework lacking temporal or source integration, underlines the model’s robust capability for CAG, effectively managing diverse flawed information and affirming the universality of CAG.
@@ 5.3 Analysis Study
In the following, we will present several analysis against the robustness and limitation of current Credibility-aware Generation.
@@ 5.3.1 Noise Robustness Analysis
Previous research has demonstrated that an increase in the proportion of noise within the context significantly degrades model performance (Xie et al., 2023; Chen et al., 2023b). To assess the robustness of diverse methods against flawed information, we vary the ratio of noisy documents within the total document set across three distinct datasets: RGB, EvolvingTempQA and NewsPollutedQA, and observe the consistency in performance changes across different models as the ratio of noisy documents increased.
We present the results in Figure 3 and can see that: Credibility-aware Generation makes the model robust to flawed information, which enhances its ability to discern and prioritize accurate information. As the proportion of noise in the context increases, most of the models exhibit performance degradation aligning with the observations made by Chen et al. (2023b). However, our models show greater robustness compared to others, notably displaying performance improvements on EvolvingTempQA as the noise ratio rises from 0.4 to 0.6.
@@ 5.3.2 Analysis of Discarding Low Credibility Documents
Upon assigning credibility to the documents in context, an alternative intuitive strategy is to simply discard low credibility documents. However, considering that credibility assessments are not precise, this strategy may inadvertently filter out helpful information, thereby impairing the accuracy of the model’s responses. To demonstrate this, we compare the performance of LLMs in this setting with that of CAG-7B in Open-domain QA. The results are shown in Figure 4, we can clearly see that: by preserving more document information and differentiating them based on explicit credibility in the prompt, our CAG framework mitigates the risk of losing valuable information. As a result, the accuracy and comprehensiveness of the responses are improved.
@@
Figure 4: The comparison of performance of LLMs under discarding low credibility document setting and CAG-7B model across different Open-domain QA datasets. We can see that the CAG-7B significantly outperforms other models under this strategy.
5.3.3 Effect of Credibility Annotation Accuracy
@@
6 Customized Credibility Applications
@@
6.1 Personalized Response Generation
@@
Figure 5: CAG provides personalized responses. We can see that CAG combines with user preferences to leverage customized credibility, offering personalized responses.
@@
6.2 Knowledge Conflict Resolution
@@
7 Related Work
Retrieval-Augmented Generation (Lewis et al., 2020) integrates a retriever with a generator to improve text generation quality by utilizing external knowledge (Izacard and Grave, 2021; Borgeaud et al., 2022; Shi et al., 2023b). However, the accuracy of RAG is compromised by flawed information, as the inclusion of noisy (Chen et al., 2023b; Kasai et al., 2022), outdated (Wang et al., 2023a), or false information (Chen and Shu, 2023; Pan et al., 2023) during the retrieval negatively impacts the generator’s outputs.
Previous studies have primarily focused on filtering, ranking, or manually evaluating retrieved documents to mitigate the impact of flawed information. For instance, Peng et al. (2023); Wang et al. (2023b) deploy various filtering algorithms to remove irrelevant text. Zhang et al. (2023) utilizes document timestamps to identify and discard outdated information. However, these approaches are limited by the accuracy of filtering algorithms, thereby discarding helpful information and impairing the effectiveness of Retrieval-Augmented Generation. Meanwhile, misinformation is primarily addressed by identifying falsehoods through fact-checking (Vijjali et al., 2020). However, this approach necessitates either human verification or further training of the discriminator Baek et al. (2023), both of which can be resource-intensive and potentially introduce bias (Draws et al., 2022; Oeldorf-Hirsch et al., 2023; Su et al., 2023). In comparison, our work mitigates the impact of flawed information without discarding documents by introducing multi-feature dimensions of external information to assess the credibility level of each document.
8 Conclusions
This paper proposes Credibility-aware Generation to address the challenge of flawed information. To equip the model with CAG capabilities, we introduce a data transformation framework aimed at generating credibility-based dataset, upon which we fine-tune the model. To effectively verify the ability of model credibility-aware generation to handle flawed information, we construct a benchmark from different real-world scenarios. Experimental results show that our model can effectively understand credibility, exhibiting robustness in the face of flawed information and significantly outperforming other models with retrieval augmentation.
Moreover, our framework is widely applicable to various real-world scenarios, offering customizable, reliable, and controllable outcomes. For instance, by constructing a unique interest library and profile for each user, and assigning credibility to retrieved documents based on this profile, personalized responses can be generated accordingly. This paper also sheds light on many future directions such resolving knowledge conflicts and designing more systems to incorporate external knowledge into LLMs.
Limitations
There are several limitations of our current CAG framework, which we plan to address in the future. Firstly, in the open-domain QA annotations, we employ automatic credibility assessment instead of relying on given golden paragraphs to better simulate real-world scenarios. Performance gaps exist due to retriever and segmentation strategy limitations. Future work will explore advanced credibility assessment algorithms to improve our model’s performance with flawed information. Secondly, our methodology, effectively applied to RAG, acknowledges the broader research domain encompassing external resources like knowledge graphs and tool usage. We aim to expand our work to domains requiring diverse external information integration, including retrieved data, knowledge graph data, and tool output.
Ethics Statement
In the following we will briefly state the moral hazard we may be involved in. Section 4.2 introduces a dataset manually labeled by members of our research team, all of whom are graduate students specializing in NLP. In Section 4.3, we examine how LLMs employ credibility processing mechanisms to address disinformation in an environment rife with false information. Our study involves experimental settings using ChatGPT to generate fake news through prompts. It is crucial to emphasize that these experiments are strictly for research purposes, do not involve any personal privacy information, and will not be used for any other purposes.
Appendix A Appendix
A.1 Statistics on training data
The composition and statistics of the training data are as follows:
Task Dataset Train (#)
Dialogue ShareGPT (Chiang et al., 2023) 3426
ODQA HotpotQA(Yang et al., 2018a) 5287
ELI5 (Fan et al., 2019) 2000
QAMPARI(Amouyal et al., 2023) 1000
WikiQA(Yang et al., 2015) 1040
MRC NewsQA(Trischler et al., 2017) 2135
PubmedQA(Jin et al., 2019) 12552
A.2 A comparison of CAGB with other similar benchmarks
@@
A.3 Prompts used on the CAGB
We conduct an evaluation of ASQA utilizing the prompts provided in Gao et al. (2023). The prompts utilized for the evaluation of the NewsPollutedQA dataset, under the settings of retrieval-based, retrieval and reranking, and retrieval and credibility in the zero-shot scenario, are displayed in Figures 7 and 8. The prompts used for assessing other datasets, under the settings of retrieval-based, retrieval and reranking, and retrieval and credibility in the zero-shot scenario, can be found in Figure 9 and Figure 10.

### v3

@@ 5 Experiments
To demonstrate the effectiveness of our framework in handling flawed information in real-world QA scenarios, we conduct comprehensive experiments within the CAGB. All these results verify the effectiveness of the CAG framework and the corresponding training algorithm. Additionally, our models maintain robustness even with the increasing noise in the context. In the following sections, we will discuss our experiments and conclusions in detail.
@@ 5.1 Setup
We compare our method with the following three strategies incorporated with 7 LLMs across various scales:
@@
We evaluate advanced models, including ChatGPT (gpt-3.5-turbo-0613), LLaMA-2-7B, 13B, 70B, Vicuna-7B-v1.5 and Mistral-7B-Instruct (Jiang et al., 2023). Additionally, we create a dataset mirroring the model training data but without credibility annotations and with initial answers, on which we fine-tune the LLaMA-2-7B model, and named the trained model vanilla IFT.
Figure 3: The performance of LLMs under varying noise ratios, which denote the proportions of retrieved noise documents. As the noise ratio increases, the performance of other methods markedly declines; in contrast, our model maintains stable performance in high noise ratio, attributed to its enhanced ability to prioritize accurate information.
@@
We use LLaMA-2-7B, 13B and Mistral-7B as our base models. We train the LLaMA-2 model and the Mistral 7B model for 3 epochs with a learning rate of 1e-5 on A100-80G GPUs.To provide relevance scores, we use SPLADE (Formal et al., 2021) as our retriever. For all language models, we include 3-shot QA examples within the prompt. We utilize Exact Match (EM) (Stelmakh et al., 2022) as the primary evaluation metric for all datasets. It is calculated by checking whether the short answers provided are exact substrings of the generation. We set the temperature to 0.01 during the inference. Additional experimental settings and the prompts used for evaluation and are provided in the Appendix A.11. These models can generate answers with reasoning processes based on the given prompt. Moreover, Appendix A.7 shows the results of the experiments using CoT prompt consistent with that used to generate training data.
@@ 5.2 Overall Results
The main results of the three scenarios are shown in the Table 2, we can clearly see that our model efficiently utilizes credibility to provide more accurate and credible responses. In the following, we analyze the experimental results in detail:
@@
In scenarios including open-domain QA, time-sensitive QA, and misinformation pollutedQA, existing LLMs, including ChatGPT and LLaMA-2-70B, face challenges due to interference from flawed information. In the retrieval-based open-domain QA, the average EM score for ChatGPT is only 41.5%, while 44.1% for LLaMA-2-70B. All models exhibit low performance on the Musique, NewsPollutedQA, which are characterized by high ratios of flawed information. Reranking with external relevance scores can assist the model to a certain extent, as the model is sensitive to the order of documents (Xie et al., 2023).
@@
The models, developed through training on LLaMA 7B, 13B, and Mistral 7B with CAG, not only exhibit improved reliability in its outputs but also excel in new, challenging situations, including time-sensitive QA and misinformation polluted QA. This performance, achieved within an open-domain QA framework lacking temporal or source integration, effectively manages diverse flawed information and affirms the universality of CAG.
@@ 5.3 Analysis Study
In the following, we will present analysis against the robustness and limitation of current CAG model. Due to the space limit, experimental results on the effect of credibility annotation accuracy are shown in the Appendix A.2.
@@ 5.3.1 Noise Robustness Analysis
Previous research has demonstrated that an increase in the proportion of noise within the context significantly degrades model performance (Xie et al., 2023; Chen et al., 2023). To assess the robustness of diverse methods against flawed information, we vary the ratio of noisy documents across three distinct datasets: RGB, EvolvingTempQA and NewsPollutedQA, and observe the consistency in performance changes across different models.
We present the results in Figure 3 and can see that: Credibility-aware Generation makes the model robust to flawed information, which enhances its ability to discern and prioritize accurate information. As the proportions of noise in the context increases, most of the models exhibit performance degradation aligning with the observations made by Chen et al. (2023). However, our models show greater robustness compared to others, notably the improved performance of CAG-13B on EvolvingTempQA. The results of the noise robustness analysis for all LLMs are shown in the Appendix A.13.
@@ 5.3.2 Analysis of Discarding Low Credibility Documents
Figure 4: The comparison of performance of LLMs under discarding low credibility document setting and CAG-7B across six datasets.
Upon assigning credibility to the documents in context, an alternative intuitive strategy is to simply discard low credibility documents. However, considering that credibility assessments are not precise, this strategy may inadvertently filter out helpful information, thereby impairing the accuracy of the model’s responses. To demonstrate this, we compare the performance of LLMs in this setting with that of CAG-7B in our CAGB. The results are shown in Figure 4, and full results are in Appendix A.8. We can clearly see that: by preserving more document information and differentiating them based on explicit credibility in the prompt, our CAG framework mitigates the risk of losing valuable information. As a result, the accuracy and comprehensiveness of the responses are improved.
6 Related Work
Retrieval-Augmented Generation (Lewis et al., 2020) integrates a retriever with a generator to improve text generation quality by utilizing external knowledge (Izacard and Grave, 2021; Borgeaud et al., 2022; Shi et al., 2023b). However, the accuracy of RAG is compromised by flawed information, as the inclusion of noisy (Chen et al., 2023), outdated (Kasai et al., 2022; Wang et al., 2023a), or false information (Chen and Shu, 2023; Pan et al., 2023) during the retrieval negatively impacts the generator’s outputs.
Previous studies have primarily focused on filtering, ranking, or manually evaluating retrieved documents to mitigate the impact of flawed information. For instance, Peng et al. (2023); Wang et al. (2023b) deploy various filtering algorithms to remove irrelevant text. Zhang and Choi (2023) utilizes timestamps to identify and discard outdated information. However, these approaches are limited by the accuracy of filtering algorithms, thereby discarding helpful information and impairing the effectiveness of RAG. Meanwhile, misinformation is primarily addressed by identifying falsehoods through fact-checking (Vijjali et al., 2020). However, this approach necessitates either human verification or further training of the discriminator Baek et al. (2023), both of which can be resource-intensive and introduce bias (Draws et al., 2022; Su et al., 2023). In comparison, our work mitigates the impact of flawed information without discarding documents by introducing multi-feature dimensions of external information to assess the credibility level of each document.
Researchers fine-tune language models to better leverage the context provided in the input. For instance, Li et al. (2023) train the model using counterfactuals and irrelevant context to prioritize context. Yoran et al. (2023) include irrelevant context in the training samples, making the model robust to irrelevant documents. Asai et al. (2023) train the model on contexts with reflective tokens, enabling it to evaluate the relevance of passages during generation. However, these approaches focus mainly on irrelevant documents. Meanwhile, the model predominantly learns implicit rules, resulting in opaqueness of the generation.
7 Conclusions
This paper proposes Credibility-aware Generation to address the challenge of flawed information. To equip the model with CAG capabilities, we introduce a data transformation framework aimed at generating credibility-based dataset, upon which we fine-tune the model. To effectively verify the ability of model Credibility-aware Generation to handle flawed information, we construct a benchmark from different real-world scenarios. Experimental results show that our model can effectively utilize credibility, exhibiting robustness in the face of flawed information and significantly outperforming other models with retrieval augmentation.
Moreover, through customizing the credibility, our approach can be applied to the real-world scenario including personalized response generation, for which we provide a detailed case study in the Appendix A.5.
Limitations
There are several limitations of our current CAG framework, which we plan to address in the future. Firstly, we have established a flexible credibility assessment mechanism, focusing more on endowing the model with the ability to generate based on credibility. However, credibility assessment is also a crucial part, and the current performance gap exists due to the retrieval strategy and influencing factors. In future research, we will delve further into credibility assessment to enhance the performance of our model. Secondly, despite our method demonstrating strong generalization capabilities, it still relies on additional training data annotation and training. In the future, we will explore how to enable existing models to perform confidence-aware generation without the need for further training. Thirdly, our methodology, effectively applied to RAG, acknowledges the broader research domain encompassing external resources like knowledge graphs and tool usage. We aim to expand our work to domains requiring diverse external information integration, including retrieved data, knowledge graph data, and tool output.
Acknowledgements
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), the Natural Science Foundation of China (No. 62122077 and 62106251), and Beijing Municipal Science and Technology Project (Nos. Z231100010323002).
Appendix A Appendix
A.1 Overview of Training Data Statistics
The composition and statistics of the training data are as follows:
Task Dataset Train (#)
Dialogue ShareGPT (Chiang et al., 2023) 3426
ODQA HotpotQA(Yang et al., 2018) 5287
ELI5 (Fan et al., 2019) 2000
QAMPARI(Amouyal et al., 2023) 1000
WikiQA(Yang et al., 2015) 1040
MRC NewsQA(Trischler et al., 2017) 2135
PubmedQA(Jin et al., 2019) 12552
A.2 Effect of Credibility Annotation Accuracy
@@
A.3 Fine-grained Credibility Analysis
To investigate the performance differences between fine-grained credibility and the three-level credibility method we currently use, we select several representative models and datasets. The fine-grained credibility employed ranges from a credibility score of 0 to 9, based on relevance. The experimental results are presented in Table 5. We can see that the use of fine-grained credibility models may lead to a decrease in performance. Fine-grained credibility demands higher accuracy in credibility classification and greater capability from the model to understand and differentiate credibility levels.
Model HotpotQA 2WikiMHQA MuSiQue EvolvingTempQA
ChatGPT 39.1 $\downarrow$3.1 36.0 $\downarrow$ 4.2 23.6 $\uparrow$5.4 66.0 $\downarrow$1.3
Vicuna-7B 27.9 $\downarrow$7 28.4 $\uparrow$ 1.8 11.4 $\uparrow$2.3 62.4 $\downarrow$1.8
LLaMA-2-7B 28.5 $\downarrow$ 9.1 26.4$\uparrow$8.8 13.4 $\downarrow$ 0.6 45.3$\downarrow$ 3.3
LLaMA-2-13B 34.1 $\downarrow$1.9 33 $\downarrow$ 5.4 15$\downarrow$1.4 49.9$\downarrow$2.1
Table 5: Performance of models using fine-grained credibility. The number following the downward arrow indicates the performance degradation compared to the currently used credibility granularity.
A.4 Retain the documents with the highest similarity
We conduct experiments on multi-hop QA datasets under the setting that only the most similar documents are retained. Based on the number of documents required to answer questions in each dataset, we retain the top 2 documents for HotpotQA and 2WikiMHQA, and the top 5 documents for the MuSiQue dataset. We compare the performance of the model under this strategy with our CAG-7B model, as shown in Table 6. The experimental results indicate that discarding the majority of low-similarity texts may enhance model performance. However, it still does not surpass our model, which retains as much information as possible while minimizing interference from irrelevant information in the presence of high credibility documents. Additionally, relying solely on low- similarity filtering is inadequate for removing outdated and false information.
Model HotpotQA 2WikiMHQA MuSiQue
ChatGPT 0.398 0.318 0.150
Vicuna-7B 0.353 0.284 0.174
LLaMA-2-13B 0.375 0.408 0.228
CAG-7B 0.509 0.578 0.340
Table 6: Performance of models under discarding most low-similarity documents.
A.5 Customized Credibility Applications
@@
A.5.1 Personalized Response Generation
@@
Figure 5: CAG provides personalized responses. We can see that CAG combines with user preferences to utilize customized credibility, offering personalized responses.
@@
A.5.2 Knowledge Conflict Resolution
@@
A.6 ASQA Full Results
Table 7 shows all results of LLMs on ASQA.
Model Length EM Rouge-L
retrieval-based
ChatGPT 0.400 0.404 0.370
LLaMA-2-7B 41.6 26.8 31.0
Vicuna-7B-v1.5 65.4 35.8 36.6
Mistral-7B-Instruct 25.7 30.0 34.0
LLaMA-2-13B 30.7 32.1 33.6
LLaMA-2-70B 16.1 31.6 31.6
vanilla IFT 23.7 15.7 23.1
retrieval and reranking
ChatGPT 40.8 40.2 36.9
LLaMA-2-7B 38.1 37.5 32.5
Vicuna-7B-v1.5 66.1 49.4 38.5
Mistral-7B-Instruct 24.5 41.4 35.7
LLaMA-2-13B 30.0 39.0 34.9
LLaMA-2-70B 16.3 38.8 33.0
vanilla IFT 23.8 17.6 23.0
retrieval and credibility
ChatGPT 30.4 44.0 38.5
LLaMA-2-7B 54.2 39.4 34.2
Vicuna-7B 64.9 49.0 38.5
Mistral-7B-Instruct 52.3 46.3 39.2
LLaMA-2-13B 39.1 38.5 33.6
LLaMA-2-70B 49.6 49.2 39.7
vanilla IFT 3.4 30.5 9.2
CAG-7B 94.0 50.3 39.3
CAG-13B 80.4 52.5 40.3
CAG-mistral-7B 69.7 50.5 40.3
Table 7: All results of LLMs on ASQA. The results of EM and Rouge-L are displayed multiplied by 100. indicates result reported from Gao et al. (2023).
A.7 Experimental Results Using CoT prompt
Figure 7 shows the CoT prompt which is consistent with the generation process of training data.
Figure 7: The CoT prompt used for evaluation.
Table 8 demonstrates the experimental results of the models using CoT prompt on CAGB.
Model HotpotQA 2WikiMHQA MuSiQue ASQA RGB EvolvingTempQA NewsPollutedQA
ChatGPT+CoT 0.46 0.398 0.262 0.448 0.883 0.806 0.450
Vicuna-7B+CoT 0.316 0.506 0.202 0.479 0.823 0.766 0.392
Mistral-7B-Instruct+CoT 0.392 0.314 0.102 0.475 0.74 0.752 0.394
LLaMA-2-13B+CoT 0.336 0.33 0.158 0.389 0.78 0.478 0.273
LLaMA-2-70B+CoT 0.382 0.432 0.254 0.480 0.837 0.58 0.298
CAG-7B (ours) 0.509 0.578 0.340 0.496 0.897 0.826 0.442
CAG-13B (ours) 0.514 0.604 0.408 0.525 0.917 0.829 0.483
Table 8: Experimental results of models using CoT prompt on the CAGB.
A.8 Full Results Under the Discarding Low Credibility Documents
Table 9 shows the full results under the discarding low credibility documents setting.
Model HotpotQA 2WikiMHQA MuSiQue ASQA RGB EvolvingTempQA NewsPollutedQA
ChatGPT 0.372 0.323 0.150 0.433 0.760 0.769 0.291
LLaMA-2-7B 0.286 0.306 0.178 0.390 0.710 0.610 0.192
Vicuna-7B 0.292 0.206 0.114 0.454 0.737 0.791 0.215
Mistral-7B-Instruct 0.366 0.288 0.146 0.468 0.757 0.800 0.223
LLaMA-2-13B 0.352 0.356 0.176 0.382 0.783 0.629 0.179
LLaMA-2-70B 0.418 0.316 0.317 0.477 0.840 0.688 0.173
vanilla IFT 0.350 0.354 0.230 0.302 0.723 0.642 0.408
CAG-7B (ours) 0.509 0.578 0.340 0.496 0.897 0.826 0.442
CAG-13B (ours) 0.514 0.604 0.408 0.525 0.917 0.829 0.483
Table 9: Full experimental results under the discarding low credibility documents setting.
A.9 Details of Credibility Assessment
The process of credibility assessment also encompasses the determination of a temporal threshold. The method we employ is designing prompts that allow the LLM to assess the timeliness of news articles regarding the question within varying temporal scopes. This approach takes into account the inherent validity period of the events within the question. In order to ensure the stability of the validity period evaluation, we conduct three trials, voting to select the validity period within each question. The prompt that we design can be found in Figure 8.
Figure 8: The prompt used to evaluate the validity period.
A.10 A Comparison of CAGB with Other Similar Benchmarks
@@
A.11 Prompts Used on the CAGB
We conduct an evaluation of ASQA utilizing the prompts provided in Gao et al. (2023). The prompts utilized for the evaluation of the NewsPollutedQA dataset, under the settings of retrieval-based, retrieval and reranking, and retrieval and credibility in the zero-shot scenario, are displayed in Figures 9 and 10. The prompts used for assessing other datasets, under the settings of retrieval-based, retrieval and reranking, and retrieval and credibility in the zero-shot scenario, can be found in Figure 11 and Figure 12.